# Batched Nonparametric Contextual Bandits

### v3

@@ A.1 Proof of Lemma 2
$\mathrm{KL}(Q_{M}^{T_{i-1}},Q_{i}^{T_{i-1}})$ $\leq\frac{1}{2}\sum_{t=1}^{T_{i-1}}\mathbb{E}_{Q_{M}}[z_{i}^{-2(\beta+d-\alpha\beta)}\sum_{j=1}^{s_{i}}\mathbf{1}\{X_{t}\in C_{i,j}\}\mathbf{1}\{\pi_{t}(X_{t})=1\}]$
$=\frac{1}{2}z_{i}^{-2(\beta+d-\alpha\beta)}\sum_{t=1}^{T_{i-1}}[\sum_{j=1}^{s_{i}}\mathbb{P}_{Q_{M}}(X_{t}\in C_{i,j},\pi_{t}(X_{t})=1)]$
$\leq\frac{1}{2}z_{i}^{-2(\beta+d-\alpha\beta)}\sum_{t=1}^{T_{i-1}}\sum_{j=1}^{s_{i}}z_{i}^{-d}=\frac{1}{2}T_{i-1}z_{i}^{-(2\beta+d+(d-\alpha\beta))},$
where the last inequality is because $\mathbb{P}_{Q_{M}}(X_{t}\in C_{i,j},\pi_{t}(X_{t})=1)=z_{i}^{-d}\mathbb{P}_{Q_{M}}(\pi_{t}(X_{t})=1\mid X_{t}\in C_{i,j})\leq z_{i}^{-d}$. Since $\alpha\beta\leq 1$,
$\mathrm{KL}(Q_{M}^{T_{i-1}},Q_{i}^{T_{i-1}})\leq\frac{1}{2}T_{i-1}z_{i}^{-(2\beta+d+(d-\alpha\beta))}\leq\frac{1}{2}T_{i-1}z_{i}^{-(2\beta+d)}.$
By Pinsker’s inequality, we can conclude
$\mathrm{TV}(Q_{M}^{T_{i-1}},Q_{i}^{T_{i-1}})\leq\sqrt{\frac{1}{2}\mathrm{KL}(Q_{M}^{T_{i-1}},Q_{i}^{T_{i-1}})}\leq\frac{1}{2}\sqrt{T_{i-1}z_{i}^{-(2\beta+d)}}.$
A.2 Proof of Lemma 3
For any $k\leq i\leq M$, we have
$|Q_{M}(A_{i})-Q_{i}(A_{i})|\overset{\mathrm{(i)}}{=}|Q_{M}^{T_{i-1}}(A_{i})-Q_{i}^{T_{i-1}}(A_{i})|\overset{\mathrm{(ii)}}{\leq}\mathrm{TV}(Q_{M}^{T_{i-1}},Q_{i}^{T_{i-1}})\overset{\mathrm{(iii)}}{\leq}\frac{1}{2(M-k+1)},$ (15)
where step (i) is because $A_{i}$ can be determined by observations up to $T_{i-1}$, step (ii) uses the definition of TV, and step (iii) applies Lemma 2 and the definition of $z_{i}$. Consequently,
$\sum_{i=k}^{M}Q_{i}(A_{i})$ $=Q_{M}(A_{M})+\sum_{i=k}^{M-1}Q_{i}(A_{i})$
$=Q_{M}(A_{M})+\sum_{i=k}^{M-1}(Q_{i}(A_{i})-Q_{M}(A_{i})+Q_{M}(A_{i}))$
$\overset{\mathrm{(iv)}}{\geq}Q_{M}(A_{M})+\sum_{i=k}^{M-1}(Q_{M}(A_{i})-\frac{1}{2(M-k+1)})\geq\sum_{i=k}^{M}Q_{M}(A_{i})-\frac{1}{2}\overset{\mathrm{(v)}}{=}\frac{1}{2},$
where step (iv) uses inequality (15), and step (v) uses the fact that $\sum_{i=k}^{M}Q_{M}(A_{i})=1$.
A.3 Proof of Lemma 4
We try to lower-bound the number of mistakes we make up to $T_{i}$. By inequality (5),
$\sup_{f\in\mathcal{C}_{z}}S_{T_{i}}(\pi,f,\frac{1}{2})$ $\geq\frac{1}{2^{s}}\sum_{j=1}^{s}\sum_{t=1}^{T_{i}}\sum_{\omega_{[-j]}\in\Omega_{s-1}}\frac{1}{z^{d}}\sum_{h\in\{\pm 1\}}\mathbb{P}_{\pi,f_{\omega_{[-j]}^{h}}}(\pi_{t}(X_{t})\neq h\mid X_{t}\in C_{j})$
$\geq\frac{1}{2^{s}}\sum_{j=1}^{s}\sum_{\omega_{[-j]}\in\Omega_{s-1}}\frac{1}{z^{d}}\sum_{t=1}^{T_{i}}\int\min\{dP_{\pi,f_{\omega_{[-j]}^{-1}}}^{t},dP_{\pi,f_{\omega_{[-j]}^{1}}}^{t}\}$
$\geq\frac{1}{2^{s}}\sum_{j=1}^{s}\sum_{\omega_{[-j]}\in\Omega_{s-1}}\frac{T_{i}}{z^{d}}\int\min\{dP_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i}},dP_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i}}\},$
where the second inequality invokes Le Cam’s method, and the last inequality holds since $\int\min\{dP,dQ\}=1-\mathrm{TV}(P,Q)$, and $\mathrm{TV}(P_{\pi,f_{\omega_{[-j]}^{-1}}}^{t},P_{\pi,f_{\omega_{[-j]}^{1}}}^{t})\leq\mathrm{TV}(P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i}},P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i}})$ for $t\leq T_{i}$. We continue the lower-bound to see that
$\sup_{f\in\mathcal{C}_{z}}S_{T_{i}}(\pi,f,\frac{1}{2})$ $\geq\frac{1}{2^{s}}\sum_{j=1}^{s}\sum_{\omega_{[-j]}\in\Omega_{s-1}}\frac{T_{i}}{z^{d}}\int\min\{dP_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i}},dP_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i}}\}$
$\geq\frac{1}{2^{s}}\sum_{j=1}^{s}\sum_{\omega_{[-j]}\in\Omega_{s-1}}\frac{T_{i}}{z^{d}}\int_{A_{i}}\min\{dP_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i}},dP_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i}}\}$
$=\frac{1}{2^{s}}\sum_{j=1}^{s}\sum_{\omega_{[-j]}\in\Omega_{s-1}}\frac{T_{i}}{z^{d}}\int_{A_{i}}\min\{dP_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},dP_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}}\},$
where the last step uses the fact that under $A_{i}$, the available observations for $\pi$ at $T_{i}$ are the same as those at $T_{i-1}$. Using properties of TV, we reach
$\sup_{f\in\mathcal{C}_{z}}S_{T_{i}}(\pi,f,\frac{1}{2})$ $\geq\frac{1}{2^{s}}\sum_{j=1}^{s}\sum_{\omega_{[-j]}\in\Omega_{s-1}}\frac{T_{i}}{z^{d}}\int_{A_{i}}\min\{dP_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},dP_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}}\}$
$\geq\frac{1}{2^{s}}\cdot\frac{T_{i}}{z^{d}}\sum_{j=1}^{s}\sum_{\omega_{[-j]}\in\Omega_{s-1}}\left(P_{\pi,f_{\omega_{[-j]}^{-1}}}(A_{i})-\frac{3}{2}\mathrm{TV}(P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}})\right)$
$\geq\frac{1}{2^{s}}\cdot\frac{T_{i}}{z^{d}}\sum_{j=1}^{s}\sum_{\omega_{[-j]}\in\Omega_{s-1}}\left(P_{\pi,f_{\omega_{[-j]}^{-1}}}(A_{i})-\frac{3}{2}\sqrt{\frac{1}{2}\mathrm{KL}(P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}})}\right),$
where the second inequality applies Lemma 6, and the third inequality is due to Pinsker’s inequality. Take $z=z_{i}$ and use Lemma 5, we have
$\sup_{f\in\mathcal{C}_{z_{i}}}S_{T_{i}}(\pi,f,\frac{1}{2})$ $\geq\frac{1}{2^{s_{i}}}\cdot\frac{T_{i}}{z_{i}^{d}}\sum_{j=1}^{s_{i}}\sum_{\omega_{[-j]}\in\Omega_{s_{i}-1}}\left(P_{\pi,i,f_{\omega_{[-j]}^{-1}}}(A_{i})-\frac{3}{2}\sqrt{z_{i}^{-(2\beta+d)}T_{i-1}}\right)$
$=\frac{1}{2^{s_{i}}}\cdot\frac{T_{i}}{z_{i}^{d}}\left(s_{i}2^{s_{i}-1}Q_{i}(A_{i})-\frac{3}{2}s_{i}2^{s_{i}-1}\sqrt{\frac{1}{36(M-k+1)^{2}}}\right)$
$\geq\frac{1}{8}T_{i}z_{i}^{-\alpha\beta}\frac{1}{M-k+1},$
where the last inequality uses the assumption $Q_{i}(A_{i})\geq 1/(2(M-k+1))$. Therefore, we arrive at
@@ A.3 Proof of Lemma 4
$\sup_{f\in\mathcal{C}_{z_{i}}}R_{T_{i}}(\pi;f)\apprge T_{i}^{-\frac{1}{\alpha}}\left[\sup_{f\in\mathcal{C}_{z_{i}}}S_{t_{i}}(\pi;f)\right]^{\frac{1+\alpha}{\alpha}}\apprge T_{i}z_{i}^{-\beta(1+\alpha)}(M-k+1)^{-\frac{1+\alpha}{\alpha}}.$
@@
=$\overset{\mathrm{(iii)}}{=}32D_{\phi}^{2}z^{-(2\beta+d)}\sum_{t=1}^{n}\mathbb{P}_{\pi,f_{\omega_{[-j]}^{-1}}}^{t}(\pi_{t}(X_{t})=1\mid X_{t}\in C_{j})$
@@
For any $i\in[M]$, one has
$\int_{A_{i}}\min\{dP_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},dP_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}}\}\geq P_{\pi,f_{\omega_{[-j]}^{-1}}}(A_{i})-\frac{3}{2}\mathrm{TV}(P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}}).$
@@
We can compute
$\int_{A_{i}}\min\{dP_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},dP_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}}\}$ $=\int_{A_{i}}\frac{dP_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}}+dP_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}}-|dP_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}}-dP_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}}|}{2}$
$\geq\frac{1}{2}(P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}}(A_{i})+P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}}(A_{i}))-\mathrm{TV}(P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}})$
$=\frac{1}{2}(2P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}}(A_{i})+P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}}(A_{i})-P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}}(A_{i}))-\mathrm{TV}(P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}})$
$\overset{\mathrm{(i)}}{\geq}P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}}(A_{i})-\frac{1}{2}\mathrm{TV}(P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}})-\mathrm{TV}(P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}})$
=$\overset{\mathrm{(ii)}}{=}P_{\pi,f_{\omega_{[-j]}^{-1}}}(A_{i})-\frac{3}{2}\mathrm{TV}(P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}}),$
where step (i) is due to $|P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}}(A_{i})-P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}}(A_{i})|\leq\mathrm{TV}(P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}},P_{\pi,f_{\omega_{[-j]}^{1}}}^{T_{i-1}})$, and step (ii) uses the fact that $P_{\pi,f_{\omega_{[-j]}^{-1}}}$ and $P_{\pi,f_{\omega_{[-j]}^{-1}}}^{T_{i-1}}$ are equivalent on $A_{i}$. ∎
@@ B.1 Two clean events
Suppose that $M\leq D_{1}\log(T)$ for some constant $D_{1}>0$. With probability at least $1-1/T$, for all $1\leq i\leq M$, and $C\in\mathcal{L}_{t_{i-1}+1}$, we have
@@
¯$\bar{\mathcal{I}}_{C}$ =$=\left\{k\in\{1,-1\}:\sup_{x\in C}f^{\star}(x)-f^{(k)}(x)\leq c_{1}|C|^{\beta}\right\},$
¯$\underline{\mathcal{I}}_{C}=\left\{k\in\{1,-1\}:\sup_{x\in C}f^{\star}(x)-f^{(k)}(x)\leq c_{0}|C|^{\beta}\right\},$
@@
¯$\underline{\mathcal{I}}_{C}\subseteq\bar{\mathcal{I}}_{C}.$
@@
In words, Lemma 8 guarantees that $\mathcal{A}_{C}$ happens with high probability if $E$ holds and $\mathcal{A}_{C^{\prime}}$ holds for all the ancestors $C\text{\textquoteright}$ ’ of $C$. See Section B.5.2 for the proof.

### v4

@@ A.2 Lower bounding the regret during the $m$-th batch
$\asymp M^{-1-\alpha\beta}T_{m}z_{m}^{-\alpha\beta}\left(\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{m})-\frac{1}{2M}\right).$
By Lemma 1, we obtain
$\sup_{f\in\mathcal{F}(\alpha,\beta)}R_{T_{m}}(\pi,f)$ $\apprge T_{m}^{-\frac{1}{\alpha}}\left[\sup_{(f,\frac{1}{2})\in\mathcal{F}(\alpha,\beta)}S_{T_{m}}(\pi;f)\right]^{\frac{1+\alpha}{\alpha}}$
@@
$\apprge M^{-\frac{(\alpha\beta+1)(1+\alpha)}{\alpha}}T_{m}z_{m}^{-\beta(1+\alpha)}\left(\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{m})-\frac{1}{2M}\right)^{\frac{1+\alpha}{\alpha}}.$ (14)
Since $\sum_{k=1}^{M}\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{k})\geq 1$, there exists some $m^{\star}\in[M]$ such that $\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{m^{\star}})\geq 1/M$. Because relation (A.2) holds for any $m\in[M]$, we reach
$\sup_{f\in\mathcal{F}(\alpha,\beta)}R_{T}(\pi,f)$ $\apprge M^{-\frac{(\alpha\beta+1)(1+\alpha)}{\alpha}}T_{m^{\star}}z_{m^{\star}}^{-\beta(1+\alpha)}\left(\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{m^{\star}})-\frac{1}{2M}\right)^{\frac{1+\alpha}{\alpha}}$
$\geq M^{-\frac{(\alpha\beta+1)(1+\alpha)}{\alpha}}T_{m^{\star}}z_{m^{\star}}^{-\beta(1+\alpha)}\left(\frac{1}{M}-\frac{1}{2M}\right)^{\frac{1+\alpha}{\alpha}}$
$\asymp M^{-\frac{(\alpha\beta+2)(1+\alpha)}{\alpha}}T_{m^{\star}}z_{m^{\star}}^{-\beta(1+\alpha)}\apprge M^{-\frac{(\alpha\beta+2)(1+\alpha)}{\alpha}}\frac{T_{m^{\star}}}{T^{\gamma}_{m^{\star}-1}}M^{-2\gamma}\apprge(\frac{1}{M})^{\tilde{D}_{2}}\cdot T^{\frac{1-\gamma}{1-\gamma^{M}}}.$
A.3 Proof of helper lemmas
@@
$\overset{\mathrm{(iii)}}{=}32D_{\phi}^{2}z^{-(2\beta+d)}\sum_{t=1}^{n}\mathbb{P}_{\pi,f_{\omega_{[-j]}^{-1}}}^{t}(\pi_{t}(X_{t})=1\mid X_{t}\in C_{j})$
@@
Fix $m\in[M]$. For any $i\in I_{m}$, one has
$\sum_{j=1}^{s_{m}}P_{X}(0<\phi(2Mz_{m}(x-q_{i,j}^{m}))\leq\delta(Mz_{m})^{\beta},X\in C_{i,j}^{m})\leq M^{-d}(1+d)\delta^{\alpha}.$
@@
By direct calculation,
$\sum_{j=1}^{s_{m}}$ $P_{X}(0<\phi(2Mz_{m}(x-q_{i,j}^{m}))\leq\delta(Mz_{m})^{\beta},X\in C_{i,j}^{m})$
$=s_{m}(Mz_{m})^{-d}\int_{[0,1]^{d}}\mathbf{1}\{\phi(x)\leq\delta(Mz_{m})^{\beta}\}\mathsf{d}x$
$\leq s_{m}(Mz_{m})^{-d}\mathbf{1}\{\delta(Mz_{m})^{\beta}>1\}+s_{m}(Mz_{m})^{-d}d\delta^{1/\beta}(Mz_{m})\mathbf{1}\{\delta(Mz_{m})^{\beta}\leq 1\}$
$\leq M^{-d}\delta^{\alpha}+dM^{-d}\delta^{\alpha}=M^{-d}(1+d)\delta^{\alpha},$
where the first inequality is due to $\int_{[0,1]^{d}}\mathbf{1}\{\phi(x)\leq\delta(Mz_{m})^{\beta}\}\mathsf{d}x=1-(1-\delta^{1/\beta}(Mz_{m}))^{d}\leq d\delta^{1/\beta}(Mz_{m})$ when $\delta(Mz_{m})^{\beta}\leq 1$; the second inequality uses the definition of the indicators and the assumption $\alpha\beta\leq 1$. ∎
Lemma 4.
Fix any $n\in[T]$ and any policy $\pi$. For any $m\in[M],i\in I_{m},j\in[s_{m}]$,
$\|\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{n}-\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{n}\|_{\mathrm{TV}}\leq\sqrt{n(Mz_{m})^{-(2\beta+d)}}.$
Proof.
It suffices to bound their KL-divergence. We can compute
$\mathrm{KL}(\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{n},\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{n})$ $\overset{\mathrm{(i)}}{\leq}8\mathbb{E}_{\pi,\omega_{i,j}^{m}=-1}[\sum_{t=1}^{n}(f_{\omega_{i,j}^{m}=-1}(X_{t})-f_{\omega_{i,j}^{m}=1}(X_{t}))^{2}\mathbf{1}\{\pi_{t}(X_{t})=1\}]$
$\overset{\mathrm{(ii)}}{\leq}32D_{\phi}^{2}(Mz_{m})^{-2\beta}\mathbb{E}_{\pi,\omega_{i,j}^{m}=-1}[\sum_{t=1}^{n}\mathbf{1}\{\pi_{t}(X_{t})=1,X_{t}\in C_{i,j}^{m}\}]$
$\overset{\mathrm{(iii)}}{=}32D_{\phi}^{2}(Mz_{m})^{-(2\beta+d)}\sum_{t=1}^{n}\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{t}(\pi_{t}(X_{t})=1\mid X_{t}\in C_{i,j}^{m})$
$\overset{\mathrm{(iv)}}{\leq}32D_{\phi}^{2}(Mz_{m})^{-(2\beta+d)}n\leq 2n(Mz_{m})^{-(2\beta+d)}.$
Here, step (i) uses the standard decomposition of KL divergence and Bernoulli reward structure; step (ii) is due to the definition of $f_{\omega}$; step (iii) uses $\mathbb{P}(X_{t}\in C_{i,j}^{m})=1/(Mz_{m})^{d}$, and step (iv) arises from $\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{t}(\pi_{t}(X_{t})=1\mid X_{t}\in C_{i,j}^{m})\leq 1$ for any $1\leq t\leq n$.
By Pinsker’s inequality,
$\|\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{n}-\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{n}\|_{\mathrm{TV}}\leq\sqrt{\frac{1}{2}\mathrm{KL}(\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{n},\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{n})}\leq\sqrt{n(Mz_{m})^{-(2\beta+d)}}.$
@@ B.1 Two clean events
Suppose that $M\leq D_{1}\log(T)$ for some constant $D_{1}>0$. With probability at least 11/T1-1/T, for all $1\leq i\leq M$, and $C\in\mathcal{L}_{t_{i-1}+1}$, we have
@@
$\bar{\mathcal{I}}_{C}$ $=\left\{k\in\{1,-1\}:\sup_{x\in C}f^{\star}(x)-f^{(k)}(x)\leq c_{1}|C|^{\beta}\right\},$
$\underline{\mathcal{I}}_{C}=\left\{k\in\{1,-1\}:\sup_{x\in C}f^{\star}(x)-f^{(k)}(x)\leq c_{0}|C|^{\beta}\right\},$
@@
$\underline{\mathcal{I}}_{C}\subseteq\bar{\mathcal{I}}_{C}.$
@@
In words, Lemma 6 guarantees that $\mathcal{A}_{C}$ happens with high probability if $E$ holds and $\mathcal{A}_{C^{\prime}}$ holds for all the ancestors $C\text{\textquoteright}$ of $C$. See Section B.5.2 for the proof.